40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1828     Entry time: Thu Nov 14 15:04:25 2019
Author: Duo 
Type: DailyProgress 
Category:  
Subject: ETMY check 

The first noisemon board was installed at the ETMY station. It was a prototype board that we brought to LLO and installed there since then. I checked the data today (11/14) and its LL channels is not working.

I checked the time series in the attachment 2 and 3. There are some problem causing the circuit to saturate.

I checked the drive signal below 10Hz in attachment 4. The drive signals across all channels are the same.

Attachment 1: Screenshot_from_2019-11-14_15-02-05.png  65 kB  | Hide | Hide all
Screenshot_from_2019-11-14_15-02-05.png
Attachment 2: Screenshot_from_2019-11-14_15-08-42.png  60 kB  | Hide | Hide all
Screenshot_from_2019-11-14_15-08-42.png
Attachment 3: Screenshot_from_2019-11-14_15-15-56.png  57 kB  | Hide | Hide all
Screenshot_from_2019-11-14_15-15-56.png
Attachment 4: Screenshot_from_2019-11-14_15-30-07.png  60 kB  | Hide | Hide all
Screenshot_from_2019-11-14_15-30-07.png
ELOG V3.1.3-